OpenStudy (anonymous):

4700=2[3960cos^-1(3960/a+3960)] solve for a HELP ME

OpenStudy (anonymous):

This?\[ 4700=2\cdot 3960\cos^{-1}\left(\frac{3960}{a+3960}\right) \]

OpenStudy (anonymous):

Yes, that is the question that I am struggling with. Here is a diagram if it helps

OpenStudy (anonymous):

Okay \[ \cos\left(\frac{4700}{2\cdot 3960}\right)=\frac{3960}{a+3960} \]

OpenStudy (anonymous):

\[ a\cos\left(\frac{4700}{2\cdot 3960}\right)+3960\cos\left(\frac{4700}{2\cdot 3960}\right)=3960 \]\[ a = \frac{3960\left(1-\cos\left(\frac{4700}{2\cdot 3960}\right)\right)}{\cos\left(\frac{4700}{2\cdot 3960}\right)} \]

OpenStudy (anonymous):

A little algebra is all you need.

OpenStudy (anonymous):

The equation was actually supposed to be \[4700=2(3960\cos^{-1} (\frac{ 3960 }{ A+3960 }))\]

OpenStudy (anonymous):

Thanks for helping me by the way

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

Since multiplication is associative, I could take away the parenthesis.

OpenStudy (anonymous):

Hold on let me solve that, im not sure my answer will be correct

OpenStudy (anonymous):

816.36?

OpenStudy (cggurumanjunath):

u are right it is 816.69 miles @nathanhome